OpenStudy (anonymous):

The mean and standard deviation height (in cms) for fifty girls in year 10 at a particular School are 164.3 and 7.49 respectively. The mean height for thirty-six boys is 173.5 cm. If 4 boys of the 36 boys in year 10, are shorter than the mean height for girls find the Standard deviation of the boys’ heights. Assume all heights are normally distribute

OpenStudy (amistre64):

4/36 = 1/9 = 0.11111; or about 11.11% if im recalling this right

OpenStudy (amistre64):

164.3 - 173.5 ------------- = zscore that has a percentage of 11.11% sd

OpenStudy (amistre64):

would need to find a ztable, or a calculator that can determine the zscore. Ill look on a table from google ... zscore for .1111 = -1.22 seems best sd for boys = (164.3-173.5)/-1.22 = abt 7.45